a'. Vector (b' vector ×c' vector)=1/(a vector ×b vector)c vector


1
Expert's answer
2021-08-11T19:21:30-0400

Triple product


"\vec a\cdot(\vec b\times\vec c)=(\vec a\times\vec b)\cdot\vec c"

Given


"\vec a\cdot(\vec b\times\vec c)=\dfrac{1}{(\vec a\times\vec b)\cdot\vec c}"

Then


"\bigg(\vec a\cdot(\vec b\times\vec c)\bigg)\bigg((\vec a\times\vec b)\cdot\vec c\bigg)=1"

"\bigg(\vec a\cdot(\vec b\times\vec c)\bigg)^2=1"

"\vec a\cdot(\vec b\times\vec c)=\pm1"